### v1

# Completeness conditions for spacetimes with low-regularity metrics

Abstract.
We extend Beem’s three completeness notions—finite compactness, timelike Cauchy completeness, and Condition A —originally defined for spacetimes, to Lorentzian length spaces and study their relationships. We prove that finite compactness implies timelike Cauchy completeness and that timelike Cauchy completeness implies Condition A for globally hyperbolic Lorentzian length spaces. Furthermore, for globally hyperbolic $C^{1}$-spacetimes under suitable assumptions ensuring good behavior of causal geodesics, we show that the three conditions are equivalent. These results can be regarded as a Hopf-Rinow type theorem for low-regularity Lorentzian geometry. The appendix presents examples of $C^{1}$-spacetimes—where geodesic uniqueness may fail— in which causal geodesics nevertheless behave well, illustrating the scope of our results.
Key words and phrases:
Lorentzian geometry, Hopf–Rinow Theorem, Lorentzian length spaces
2020 Mathematics Subject Classification:
53C50, 53C23, 53B30, 51K10
1. Introduction
In Riemannian geometry, a fundamental result is that every connected Riemannian manifold is naturally endowed with a metric space structure. The Hopf–Rinow theorem asserts that several completeness conditions are equivalent in this setting.
Fact 1.1 (Hopf–Rinow, cf. [dC92, Chapter 7, Theorem 2.8]).
Let $(M,g)$ be a connected Riemannian manifold. The following conditions are equivalent:
  1. (1)
    $M$ is boundedly compact, i.e., every closed metric ball in $M$ is compact;
  2. (2)
    $M$ is metrically complete, i.e., every Cauchy sequence converges in $M$;
  3. (3)
    $M$ is geodesically complete, i.e., for all $p\in M$, any geodesic $\gamma\colon I\to M$ starting at $p$ is defined for all $t\in\mathbf{R}$.
This theorem can be extended to locally compact length spaces with slight modifications.
Fact 1.2 (Hopf, Rinow, Cohn-Vossen, cf. [BBI01, Theorem 2.5.28]).
Let $X$ be a locally compact length space. The following conditions are equivalent:
  1. (1)
    $X$ is boundedly compact;
  2. (2)
    $X$ is metrically complete;
  3. (3)
    Every geodesic $\gamma\colon[0,a)\to X$ can be extended to a continuous path $\tilde{\gamma}\colon[0,a]\to X$.
In contrast with the Riemannian case, the completeness conditions for Lorentzian manifolds are considerably more involved. In particular, for Lorentzian manifolds with low-regularity metrics, much less is known, and many fundamental questions remain open.
Beem [B76] introduced finite compactness and timelike Cauchy completeness, which are Lorentzian analogues of bounded compactness and metric completeness for Riemannian manifolds, respectively. Beem proved a following Lorentzian version of the Hopf–Rinow theorem for $C^{2}$-spacetimes. Here, a $C^{2}$-spacetime means a smooth manifold endowed with a Lorentzian metric of class $C^{2}$.
Fact 1.3 (Beem [B76, Theorem 5]).
Let $(M,g)$ be a globally hyperbolic $C^{2}$-spacetime. The following conditions are equivalent:
  1. (1)
    $M$ is finitely compact (cf. Definition 3.1);
  2. (2)
    $M$ is timelike Cauchy complete (cf. Definition 3.3);
  3. (3)
    $M$ satisfies Condition A (cf. Definition 3.7).
A recent result by Burtscher and García-Heveling [BG24] provides a new Lorentzian version of the Hopf–Rinow theorem, characterizing global hyperbolicity via the null distance introduced by Sormani and Vega [SV16], though its current formulation appears to require at least $C^{2,1}$ regularity.
Therefore, the question of completeness remains poorly understood for spacetimes with Lorentzian metrics of low-regularity.
In recent years, there has been significant progress on low-regularity Lorentzian geometry via a synthetic approach. A central framework in this direction is provided by Lorentzian length spaces, introduced by Kunzinger and Sämann [KS18]. These spaces admit a causal structure and a time separation function without assuming a manifold structure or even a Lorentzian metric. They provide a general framework extending classical Lorentzian geometry, allowing one to study causal and metric properties in low-regularity settings. For example, Beran, Ohanyan, Rott, and Solis [BORS23] establish a Lorentzian splitting theorem in this setting.
It is thus natural to ask how far Beem’s Hopf–Rinow-type implications extend to spacetimes with low-regularity Lorentzian metrics and, more broadly, to Lorentzian length spaces. Furthermore, within this synthetic framework, we may ask to what extent one can verify which classical completeness implications persist—thereby assessing how well Lorentzian length spaces provide a suitable framework.
In this article, we extend Beem’s completeness conditions [B76] to Lorentzian length spaces and show that two of Beem’s implications continue to hold in this general setting.
Theorem 3.9.
Let $(X,\ll,\leq,d,\tau)$ be a globally hyperbolic Lorentzian length space.
  1. (1)
    If $(X,\ll,\leq,d,\tau)$ is finitely compact, then $(X,\ll,\leq,d,\tau)$ is timelike Cauchy complete.
  2. (2)
    If $(X,\ll,\leq,d,\tau)$ is timelike Cauchy complete, then $(X,\ll,\leq,d,\tau)$ satisfies Condition A.
We also show that the remaining implication holds for $C^{1}$-spacetimes under appropriate geodesic assumptions. Combined with Theorem 3.9, this yields the following result.
Theorem 4.12.
Let $(M,g)$ be a connected, globally hyperbolic, causally non-branching, and proper $C^{1}$-spacetime such that the exponential map as defined in Definition 4.8 is continuous. Then the following three conditions are equivalent:
  1. (1)
    $M$ is finitely compact;
  2. (2)
    $M$ is timelike Cauchy complete;
  3. (3)
    $M$ satisfies Condition A.
The article is organized as follows. In Section 2, we review the notion of Lorentzian length spaces introduced by Kunzinger and Sämann [KS18] and collect some results to be used in the proof of the main theorems. In Section 3, we extend the three completeness conditions to Lorentzian length spaces and establish Theorem 3.9. In Section 4, we study geodesics in $C^{1}$-spacetimes, analyze the Lorentzian distance along causal geodesics, and establish Theorem 4.12. In Appendix A, we present examples of spacetimes with Lorentzian metrics of regularity below $C^{1,1}$ for which geodesic uniqueness still holds, even though classical uniqueness results for geodesics generally fail in this regularity class. In particular, we exhibit $C^{1}$-spacetimes that are not $C^{1,1}$ and that satisfy the assumptions of Theorem 4.12. These spacetimes lie outside the scope of Fact 1.3 but are covered by Theorem 4.12.
Acknowledgements.
The author would like to thank Kotaro Yamada for helpful discussions and valuable advice. This work was supported by JST SPRING, Japan Grant Number JPMJSP2180.
2. Preliminaries
In this section, we recall some notions and results on Lorentzian (pre-)length spaces, which will be used in lemmas and propositions leading to the main theorems. Throughout the paper, we denote by $\mathbf{N}$ the set of natural numbers starting from $1$.
Definition 2.1 ([KS18, Definition 2.1]).
A causal space $(X,\ll,\leq)$ is a set $X$ together with two binary relations $\ll$ and $\leq$ where both are transitive, $\leq$ is reflexive, and all $x,y\in X$ with $x\ll y$ satisfy $x\leq y.$ We write $x<y$ if $x\leq y$ and $x\neq y$.
The chronological future $I^{+}(x)$, chronological past $I^{-}(x)$, causal future $J^{+}(x)$, and causal past $J^{-}(x)$ of $x\in X$ are defined by
(2.1) $I^{+}(x)$ $\coloneqq\{y\in X;x\ll y\},\ I^{-}(x)\coloneqq\{y\in X;y\ll x\},$
(2.2) $J^{+}(x)$ $\coloneqq\{y\in X;x\leq y\},\ J^{-}(x)\coloneqq\{y\in X;y\leq x\}.$
Definition 2.2 ([KS18, Definition 2.8]).
A Lorentzian pre-length space $(X,\ll,\leq,\penalty-10000d,\tau)$ is a causal space $(X,\ll,\leq)$ together with a distance $d$ on $X$ and a map $\tau\colon X\times X\to[0,\infty]$ which satisfies the following properties:
  1. (i)
    $\tau$ is lower semicontinuous (with respect to the topology induced by $d$),
  2. (ii)
    $\tau(x,z)\geq\tau(x,y)+\tau(y,z)$ for all $x,y,z\in X$ with $x\leq y\leq z,$
  3. (iii)
    $\tau(x,y)>0$ if and only if $x\ll y,$
  4. (iv)
    $\tau(x,y)=0$ if $x\not\leq y.$
$\tau$ is called the time separation function.
Remark 2.3.
In the following sections, we always use the metric topology induced by $d$ for all Lorentzian (pre-)length spaces.
Fact 2.4 ([KS18, Lemma 2.10]).
Let $(X,\ll,\leq,d,\tau)$ be a Lorentzian pre-length space and let $x,y,z\in X$ with $x\leq y\ll z$ or $x\ll y\leq z.$ Then $x\ll z.$
Definition 2.5 ([KS18, Definition 2.18]).
Let $(X,\ll,\leq,d,\tau)$ be a Lorentzian pre-length space and $I\subset\mathbf{R}$ an interval. A future-directed causal (resp. timelike) curve is a non-constant, locally Lipschitz continuous map $\gamma\colon I\to X$ satisfying $\gamma(t_{1})\leq\gamma(t_{2})\ (\text{resp.}\ \gamma(t_{1})\ll\gamma(t_{2}))$ for all $t_{1},t_{2}\in I$ with $t_{1}<t_{2}$. A past-directed causal (resp. timelike) curve is a non-constant, locally Lipschitz continuous map $\gamma\colon I\to X$ satisfying $\gamma(t_{2})\leq\gamma(t_{1})\ (\text{resp.}\ \gamma(t_{2})\ll\gamma(t_{1}))$ for all $t_{1},t_{2}\in I$ with $t_{1}<t_{2}$.
We introduce the length of causal curves by using the time separation function.
Definition 2.6 ([KS18, Definition 2.24]).
Let $(X,\ll,\leq,d,\tau)$ be a Lorentzian pre-length space and $\gamma\colon[a,b]\to X$ a future-directed causal curve. The $\tau$-length of $\gamma$ is defined by
(2.3) $L_{\tau}(\gamma)\coloneqq\inf\left\{\sum_{i=0}^{n-1}\tau(\gamma(t_{i}),\gamma(t_{i+1}));n\in\mathbf{N},a=t_{0}<t_{1}<\dots<t_{n}=b\right\}.$
If $\gamma$ is a past-directed causal curve, the $\tau$-length of $\gamma$ is defined by
(2.4) $L_{\tau}(\gamma)\coloneqq\inf\left\{\sum_{i=0}^{n-1}\tau(\gamma(t_{i+1}),\gamma(t_{i}));n\in\mathbf{N},a=t_{0}<t_{1}<\dots<t_{n}=b\right\}.$
Definition 2.7 ([KS18, Definition 2.33]).
Let $(X,\ll,\leq,d,\tau)$ be a Lorentzian pre-length space. A future-directed causal curve $\gamma\colon[a,b]\to X$ is called maximal if $L_{\tau}(\gamma)=\tau(\gamma(a),\gamma(b))$.
To avoid some technical difficulties, we impose the following three conditions, which are used in the definition of Lorentzian length spaces.
Definition 2.8 ([KS18, Definition 3.1]).
A Lorentzian pre-length space $(X,\ll,\leq,\penalty-10000d,\tau)$ is called causally path connected if for all $x,y\in X$ with $x\ll y$, there exists a future-directed timelike curve from $x$ to $y$, and for all $x<y$, there exists a future-directed causal curve from $x$ to $y.$
Definition 2.9 ([KS18, Definition 3.4]).
Let $(X,\ll,\leq,d,\tau)$ be a Lorentzian pre-length space and $x\in X.$ A neighborhood $U$ of $x$ is called causally closed if for any sequences $(p_{n})_{n},(q_{n})_{n}$ in $U$ satisfying $p_{n}\leq q_{n}$ for all $n\in\mathbf{N}$ and $p_{n}\to p\in\overline{U},\ q_{n}\to q\in\overline{U}$, then $p\leq q$, where $\overline{U}$ is the closure of $U.$ A Lorentzian pre-length space $(X,\ll,\leq,d,\tau)$ is called locally causally closed if each point in $X$ has a causally closed neighborhood.
Definition 2.10 ([KS18, Definition 3.16]).
A Lorentzian pre-length space $(X,\ll,\penalty-10000\leq,d,\tau)$ is called localizable if for all $x\in X,$ there exists an open neighborhood $\Omega_{x}\subset X$ of $x$ which satisfies the following properties:
  1. (i)
    there is a constant $C>0$ such that the $d$-length of $\gamma$ given by
    (2.5) $L^{d}(\gamma)\coloneqq\sup\left\{\sum_{i=0}^{n-1}d(\gamma(t_{i}),\gamma(t_{i+1}));n\in\mathbf{N},a=t_{0}<t_{1}<\dots<t_{n}=b\right\}$
    satisfies $L^{d}(\gamma)\leq C$ for all causal curves $\gamma$ contained in $\Omega_{x};$
  2. (ii)
    there exists a continuous map $\omega_{x}\colon\Omega_{x}\times\Omega_{x}\to[0,\infty)$ which satisfies the following properties:
    1. (a)
      $(\Omega_{x},\ll|_{\Omega_{x}\times\Omega_{x}},\leq|_{\Omega_{x}\times\Omega_{x}},d|_{\Omega_{x}\times\Omega_{x}},\omega_{x})$ is a Lorentzian pre-length space,
    2. (b)
      $I^{\pm}(y)\cap\Omega_{x}\neq\emptyset$ for every $y\in\Omega_{x}$;
  3. (iii)
    for all $p,q\in\Omega_{x}$ with $p<q$, there exists a future-directed causal curve $\gamma_{p,q}$ from $p$ to $q$ which satisfies the following properties:
    1. (a)
      $L_{\tau}(\gamma_{p,q})\geq L_{\tau}(\lambda)$ for all future-directed causal curves $\lambda$ from $p$ to $q$ contained in $\Omega_{x},$
    2. (b)
      $L_{\tau}(\gamma_{p,q})=\omega_{x}(p,q)\leq\tau(p,q).$
If every point $x\in X$ has a neighborhood basis of open sets $\Omega_{x}$ satisfying (i)-(iii), then $(X,\ll,\leq,d,\tau)$ is called strongly localizable.
We now define the main object: Lorentzian length spaces.
Definition 2.11 ([KS18, Definition 3.22]).
A Lorentzian length space is a causally path connected, locally causally closed and localizable Lorentzian pre-length space satisfying
(2.6) $\tau(x,y)=\begin{cases}\displaystyle\sup_{\gamma\in\Omega_{x,y}}L_{\tau}(\gamma),&\Omega_{x,y}\neq\emptyset,\\ 0,&\Omega_{x,y}=\emptyset,\end{cases}$
where $\Omega_{x,y}$ is the set of all future-directed causal curves from $x$ to $y$.
A fundamental source of Lorentzian length spaces comes from spacetimes with low-regularity Lorentzian metrics. However, one cannot go below Lipschitz regularity in general: for metrics of regularity strictly less than $C^{0,1}$, there exist spacetimes that fail even to be Lorentzian pre-length spaces (cf. [KS18, Example 5.2]).
In contrast, for spacetimes with continuous metrics satisfying suitable assumptions, the standard construction yields Lorentzian length spaces. Fix any complete Riemannian metric $h$ on $M$ and let $d^{h}$ be the associated Riemannian distance. Let $T$ be the usual Lorentzian distance induced by $g$ (cf. Definition 4.3).
Fact 2.12 ([KS18, Proposition 5.12]).
Let $(M,g)$ be a $C^{0}$-spacetime that is causally plain (cf. [CG12, Definition 1.16]) and strongly causal (cf. [O83, Definition 14.11]). Then, $(M,\ll,\leq,d^{h},T)$ is a strongly localizable Lorentzian length space. In particular, if $(M,g)$ is a strongly causal $C^{0,1}$-spacetime, then $(M,\ll,\leq,d^{h},T)$ is a strongly localizable Lorentzian length space.
To avoid pathological examples, we need to impose some causal conditions. We now introduce the following conditions.
Definition 2.13 ([KS18, Definition 2.35]).
A Lorentzian length space $(X,\ll,\leq,\penalty-10000d,\tau)$ is called
  1. (1)
    causal if $x\leq y$ and $y\leq x$ implies $x=y$ for all $x,y\in X$,
  2. (2)
    non-totally imprisoning if for every compact set $K\subset X$, there is $C>0$ such that the $d$-length in (2.5) of all causal curves in $K$ is bounded by $C$,
  3. (3)
    globally hyperbolic if $(X,\ll,\leq,d,\tau)$ is non-totally imprisoning and the causal diamond $J(x,y)\coloneqq J^{+}(x)\cap J^{-}(y)$ is compact for all $x,y\in X.$
We collect basic consequences of global hyperbolicity that we shall use repeatedly.
Fact 2.14 ([KS18, Theorem 3.26 (ii)]).
Let $(X,\ll,\leq,d,\tau)$ be a Lorentzian length space. If $(X,\ll,\leq,d,\tau)$ is non-totally imprisoning, then $(X,\ll,\leq,d,\tau)$ is causal.
Fact 2.15 ([ACS20, Proposition 3.13]).
Let $(X,\ll,\leq,d,\tau)$ be a globally hyperbolic Lorentzian length space. Then the sets $J^{\pm}(x)$ are closed for all $x\in X$.
Fact 2.16 ([KS18, Theorem 3.28]).
Let $(X,\ll,\leq,d,\tau)$ be a globally hyperbolic Lorentzian length space. Then $\tau$ is (i) finite, i.e., $\tau(x,y)\neq\infty$ for all $x,y\in X$, and (ii) continuous.
Fact 2.17 ([KS18, Theorem 3.30]).
Let $(X,\ll,\leq,d,\tau)$ be a globally hyperbolic Lorentzian length space. Then $(X,\ll,\leq,d,\tau)$ is geodesic, i.e., for all $x,y\in X$ with $x<y$, there exists a maximal future-directed causal curve $\gamma$ from $x$ to $y$.
3. Completeness conditions and their implications
In this section, we extend three conditions—finite compactness, timelike Cauchy completeness, and Condition A—to Lorentzian pre-length spaces, and prove Theorem 3.9. Theorem 3.9 follows directly from Theorem 3.4 and Theorem 3.8.
The following definition is a Lorentzian analogue of bounded compactness. Beem [B76] first introduced it for spacetimes.
Definition 3.1.
A Lorentzian pre-length space $(X,\ll,\leq,d,\tau)$ is called finitely compact if for every $B>0$, every $p,q\in X,$ and every sequence of points $(x_{n})_{n}$ in $X$ satisfying $p\ll q\leq x_{n}\ (\text{resp.}\ x_{n}\leq q\ll p)$ and $\tau(p,x_{n})\leq B\ (\text{resp.}\ \tau(x_{n},p)\leq B)$, there is an accumulation point of $(x_{n})_{n}$ in $X$.
The following lemma is a characterization of finite compactness for globally hyperbolic Lorentzian length spaces.
Lemma 3.2.
Let $(X,\ll,\leq,d,\tau)$ be a globally hyperbolic Lorentzian length space. Then $(X,\ll,\leq,d,\tau)$ is finitely compact if and only if for every $B>0$ and every $p,q\in X$ with $p\ll q,$ the set $K_{1}\coloneqq\{x\in X;p\ll q\leq x,\tau(p,x)\leq B\}$ is compact, and for every $B>0$ and every $p,q\in X$ with $q\ll p,$ the set $K_{2}\coloneqq\{x\in X;x\leq q\ll p,\tau(x,p)\leq B\}$ is compact.
Proof.
Assume $(X,\ll,\leq,d,\tau)$ is finitely compact. Let $p,q\in X$ with $p\ll q$ and let $K_{1}$ be as above. We first show that the set $K_{1}$ is closed. We can write $K_{1}=J^{+}(q)\cap\{x\in X;\tau(p,x)\leq B\}.$ By Fact 2.15, $J^{+}(q)$ is closed. From Fact 2.16 (ii), $\{x\in X;\tau(p,x)\leq B\}$ is also closed. Thus, $K_{1}$ is closed.
Let $A$ be any infinite subset of $K_{1}$. Then there is a sequence $(x_{n})_{n}$ in $A$ with pairwise distinct terms. The finite compactness of $X$ shows that the sequence $(x_{n})_{n}$ has an accumulation point in $K_{1}$, and this point is also an accumulation point of $A$. Thus, $K_{1}$ is limit-point compact (cf. [M75, Chapter 3-7]). By [M75, Theorem 7.4], $K_{1}$ is compact. Similarly, $K_{2}$ is compact.
Conversely, assume both $K_{1}$ and $K_{2}$ are compact. Let $B>0,\ p,q\in X,$ and $(x_{n})_{n}$ a sequence of points in $X$ satisfying $p\ll q\leq x_{n}$ and $\tau(p,x_{n})\leq B.$ Then by the assumption, it follows that there exists a subsequence of $(x_{n})_{n}$ which converges to $x_{\infty}\in K_{1}.$ Thus, $x_{\infty}$ is an accumulation point of $(x_{n})_{n}.$ Similarly, for the sequence of points $(x_{n})_{n}$ with $x_{n}\leq q\ll p$ and $\tau(x_{n},p)\leq B$, we can find an accumulation point of $(x_{n})_{n}.$
The following definition is a Lorentzian analogue of metric completeness. Beem [B76] first introduced it for spacetimes.
Definition 3.3.
A Lorentzian pre-length space $(X,\ll,\leq,d,\tau)$ is called timelike Cauchy complete if for any sequence of points $(x_{n})_{n}$ in $X$ and any sequence of non-negative numbers $(B_{n})_{n}$ which satisfy $x_{n}\ll x_{n+1}\ (\text{resp.}\ x_{n+1}\ll x_{n})$ for all $n\in\mathbf{N},$ $\tau(x_{n},x_{n+m})\leq B_{n}\ (\text{resp.}\ \tau(x_{n+m},x_{n})\leq B_{n})$ for all $n,m\in\mathbf{N}$ and $B_{n}\to 0$ as $n\to\infty$, the sequence $(x_{n})_{n}$ converges in $X$.
Beem [B76] first proved the following result for $C^{2}$-spacetimes. We extend this result to globally hyperbolic Lorentzian length spaces.
Theorem 3.4.
Let $(X,\ll,\leq,d,\tau)$ be a globally hyperbolic Lorentzian length space. If $(X,\ll,\leq,d,\tau)$ is finitely compact, then $(X,\ll,\leq,d,\tau)$ is timelike Cauchy complete.
Proof.
Assume $(X,\ll,\leq,d,\tau)$ is finitely compact. Let $(x_{n})_{n}$ be a sequence of points in $X$ satisfying $x_{n}\ll x_{n+1}$ and $\tau(x_{n},x_{n+m})\leq B_{n}$ where $B_{n}\to 0.$ Define $B\coloneqq\sup_{n\in\mathbf{N}}B_{n}$. Then $B$ is finite since $B_{n}\to 0.$ By setting $p\coloneqq x_{1}$ and $q\coloneqq x_{2}$, the finite compactness of $X$ shows that there is a point $x_{\infty}\in X$ such that $x_{n(k)}\to x_{\infty}.$
In order to prove the original sequence $(x_{n})_{n}$ converges to $x_{\infty}$, we claim the following:
(3.1) $\bigcap_{k\in\mathbf{N}}(J^{+}(x_{n(k)}))\cap J^{-}(x_{\infty})=\{x_{\infty}\}.$
In fact, (3.1) can be proved as follows. First, let $y\in\bigcap_{k\in\mathbf{N}}(J^{+}(x_{n(k)}))\cap J^{-}(x_{\infty}),$ then we see that $x_{n(k)}\in J^{-}(y)$ for all $k\in\mathbf{N}$ and $y\leq x_{\infty}.$ By Fact 2.15, we see that $x_{\infty}\in J^{-}(y)$, which implies $x_{\infty}\leq y.$ Thus, we obtain $x_{\infty}\leq y\leq x_{\infty}.$ From Fact 2.14, we see that $y=x_{\infty}$. Therefore, $\bigcap_{k\in\mathbf{N}}(J^{+}(x_{n(k)}))\cap J^{-}(x_{\infty})\subset\{x_{\infty}\}$.
For the reverse inclusion, it follows from the reflexivity of $\leq$ that $x_{\infty}\in J^{-}(x_{\infty})$. In order to show that $x_{\infty}\in\bigcap_{k\in\mathbf{N}}J^{+}(x_{n(k)})$, suppose that there exists $K\in\mathbf{N}$ such that $x_{\infty}\not\in J^{+}(x_{n(K)})$. Then, since $X$ is a regular topological space, there exist open sets $V$ and $W$ satisfying $J^{+}(x_{n(K)})\subset V,\ x_{\infty}\in W$, and $V\cap W=\emptyset$. Note that for all $k\geq K$, $J^{+}(x_{n(k)})\subset J^{+}(x_{n(K)})$ holds since $x_{n}\ll x_{n+1}$ and $x_{n(k)}\in J^{+}(x_{n(k)})\subset V$ for all $k\geq K$, so $(x_{n(k)})_{k}$ cannot converge to $x_{\infty}\in W$, a contradiction. This completes the proof of (3.1).
From (3.1), we see that for any open neighborhood $U$ of $x_{\infty}$, there exists $K\in\mathbf{N}$ such that $J^{+}(x_{n(K)})\cap J^{-}(x_{\infty})\subset U$. Indeed, suppose that for all $k\in\mathbf{N}$, $F_{k}\coloneqq(J^{+}(x_{n(k)})\cap J^{-}(x_{\infty}))\setminus U\neq\emptyset$. Each $J^{+}(x_{n(k)})\cap J^{-}(x_{\infty})$ is compact by global hyperbolicity, hence $F_{k}$, being a closed subset thereof, is also compact. Thus, $(F_{k})_{k}$ is a decreasing sequence of closed compact non-empty sets. By Cantor’s intersection theorem, $\cap_{k\in\mathbf{N}}F_{k}\neq\emptyset$. This contradicts (3.1) since $x_{\infty}\not\in F_{k}$. This implies that the original sequence $(x_{n})_{n}$ converges to $x_{\infty}$. ∎
We now introduce the notion of inextendible causal curves and geodesics to define Condition A (Definition 3.7).
Definition 3.5 ([KS18, Definition 3.10]).
Let $(X,\ll,\leq,d,\tau)$ be a Lorentzian pre-length space and let $\gamma\colon[a,b)\to X\ (-\infty<a<b\leq\infty)$ be a future-directed causal (resp. timelike) curve. The curve $\gamma$ is called future extendible if there exists a future-directed causal (resp. timelike) curve $\tilde{\gamma}\colon[a,b]\to X$ such that $\tilde{\gamma}|_{[a,b)}=\gamma.$ The curve $\gamma$ is called future inextendible if it is not future extendible. Similarly, we define the notion of past (in)extendibility.
Definition 3.6 ([BS23, Definition 1.7]).
Let $(X,\ll,\leq,d,\tau)$ be a Lorentzian pre-length space and $I\subset\mathbf{R}$ an interval. A future-directed causal curve $\gamma\colon I\to X$ is called geodesic if for each $t\in I,$ there exists a neighborhood $a,b]\subset I$ of $t$ such that $\gamma|_{[a,b]}$ is maximal (cf. Definition 2.7).
The following definition was first introduced by Beem [B76] for Lorentzian manifolds. We extend this condition to Lorentzian length spaces.
Definition 3.7.
Let $c\in[0,\infty].$ The Lorentzian pre-length space $(X,\ll,\leq,d,\tau)$ is said to satisfy Condition A if for any $x,y\in X$ with $x\ll y\ (\text{resp.}\ y\ll x$) and any future (resp. past) inextendible future-directed (resp. past-directed) causal geodesic $\gamma(t)$ starting at $y$ defined on $[0,c),$ it holds that $\tau(x,\gamma(t))\to\infty\ (\text{resp.}\ \tau(\gamma(t),x)\to\infty)$ as $t\to c.$
Beem [B76] first proved the following result for $C^{2}$-spacetimes. We extend this result to globally hyperbolic Lorentzian length spaces.
Theorem 3.8.
Let $(X,\ll,\leq,d,\tau)$ be a globally hyperbolic Lorentzian length space. If $(X,\ll,\leq,d,\tau)$ is timelike Cauchy complete, then $(X,\ll,\leq,d,\tau)$ satisfies Condition A.
Proof.
Let $x,y\in X$ with $x\ll y$ and let $\gamma(t)$ be a future inextendible, future-directed causal geodesic starting at $y$, defined on $[0,c)$. We argue by contradiction: if Condition A fails, we construct a sequence as in Definition 3.3 that forces $\gamma$ to be extendible.
Suppose that there exists $B>0$ such that $\tau(x,\gamma(t))\leq B$ for all $t\in[0,c).$ Let $(t_{n})_{n}$ be an increasing sequence of positive numbers such that $t_{n}\to c$ as $n\to\infty.$ Then it follows from Fact 2.4 and Fact 2.17 that there exist positive numbers $(c_{n})_{n}$ and maximal future-directed causal curves $(G_{n}(s))_{n}$ defined on $0,c_{n}]$ such that $G_{n}(0)=x$ and $G_{n}(c_{n})=\gamma(t_{n})$.
We will construct a timelike Cauchy sequence $(z_{n})_{n}.$ By the continuity of $d$, we can take $z_{1}\in G_{1}((0,c_{1}))$ satisfying $d(z_{1},\gamma(t_{1}))<1.$ We will show that if $z_{n}$ is taken, we can take $z_{n+1}$ in $G_{n+1}((0,c_{n+1}))$ satisfying $d(z_{n+1},\gamma(t_{n+1}))<\frac{1}{n+1}$ and $z_{n}\ll z_{n+1}$. We only need to show that $z_{n+1}$ can be taken satisfying $z_{n}\ll z_{n+1}$.
First, we show that $z_{n}$ can be taken satisfying $z_{n}\ll\gamma(t_{n}).$ Suppose that $G_{n}(s)\not\ll\gamma(t_{n})$ for all $s\in(0,c_{n}).$ Then, by Fact 2.16 (i) and Definition 2.2 (iii), we have $\tau(G_{n}(s),\gamma(t_{n}))=0$ for all $s\in(0,c_{n}).$ By Fact 2.16 (ii), we obtain
(3.2) $0=\lim_{s\to+0}\tau(G_{n}(s),\gamma(t_{n}))=\tau(x,\gamma(t_{n})).$
Thus, we see that $x\not\ll\gamma(t_{n}).$ However, since $x\ll y\leq\gamma(t_{n}),$ by Fact 2.4, we obtain $x\ll\gamma(t_{n})$, a contradiction. Thus we can take $z_{n}$ satisfying $z_{n}\ll\gamma(t_{n}).$
Next, we show that $z_{n+1}$ can be taken satisfying $z_{n}\ll z_{n+1}$. Suppose that $z_{n}\not\ll G_{n+1}(s)$ for all $s\in(0,c_{n+1}).$ Then, by Fact 2.16 (i) and Definition 2.2 (iii), we have $\tau(z_{n},G_{n+1}(s))=0$ for all $s\in(0,c_{n+1}).$ Thus, by Definition 2.2 (ii) and Fact 2.16 (ii), we obtain
(3.3) $0$ $\leq\tau(z_{n},\gamma(t_{n}))+\tau(\gamma(t_{n}),\gamma(t_{n+1}))$
(3.4) $\leq\tau(z_{n},\gamma(t_{n+1}))=\lim_{s\to c_{n+1}}\tau(z_{n},G_{n+1}(s))=0,$
which implies $\tau(z_{n},\gamma(t_{n}))=0.$ This contradicts $z_{n}\ll\gamma(t_{n})$ and thus we can take $z_{n+1}$ satisfying $z_{n}\ll z_{n+1}$.
Since $x\ll z_{n}\ll z_{n+1}$, we see that the sequence $(\tau(x,z_{n}))_{n}$ is increasing. By the assumption of contradiction that $(\tau(x,z_{n}))_{n}$ is bounded above by $B$, it follows that $(\tau(x,z_{n}))_{n}$ converges to some non-negative constant $b.$ We obtain
$\tau(z_{n},z_{n+m})$ $\leq\tau(x,z_{n+m})-\tau(x,z_{n})$
$\leq b-\tau(x,z_{n})\to 0\ (n\to\infty).$
Thus, by the timelike Cauchy completeness, the sequence $(z_{n})_{n}$ converges to some $z_{\infty}\in X.$ Since $d(z_{n},\gamma(t_{n}))\leq\frac{1}{n}\to 0$ as $n\to\infty,$ we see that $\gamma(t_{n})\to z_{\infty}.$ This contradicts the future inextendibility of $\gamma$. ∎
The following theorem is obtained by combining Theorem 3.4 and Theorem 3.8.
Theorem 3.9.
Let $(X,\ll,\leq,d,\tau)$ be a globally hyperbolic Lorentzian length space.
  1. (1)
    If $(X,\ll,\leq,d,\tau)$ is finitely compact, then $(X,\ll,\leq,d,\tau)$ is timelike Cauchy complete.
  2. (2)
    If $(X,\ll,\leq,d,\tau)$ is timelike Cauchy complete, then $(X,\ll,\leq,d,\tau)$ satisfies Condition A.
As shown above, in globally hyperbolic Lorentzian length spaces, finite compactness implies timelike Cauchy completeness, and timelike Cauchy completeness implies Condition A.
However, to derive finite compactness from Condition A, one needs to analyze in detail the behavior of geodesics carefully. For instance, there exists a spacetime with a $C^{1}$ Lorentzian metric, which is an example of a Lorentzian length space, in which geodesic uniqueness fails [KOSS22, Example 3.2]. In what follows, we focus on the particular class of Lorentzian length spaces, namely $C^{1}$-spacetimes, and examine additional assumptions under which finite compactness can be recovered from Condition A.
4. Geodesics in $C^{1}$-spacetimes
In this section, we investigate additional assumptions under which finite compactness can be derived from Condition A in $C^{1}$-spacetimes, which form a particular instance of Lorentzian length spaces.
To this end, we first analyze the properties of geodesics in $C^{1}$-spacetimes. In a $C^{1}$-spacetime, the Christoffel symbols are continuous, and hence geodesics can be defined in the same way as in the smooth case (cf. [G20, Definition 2.1]).
Definition 4.1.
Let $(M,g)$ be an $n$-dimensional $C^{1}$-spacetime. A $C^{2}$-curve $\gamma\colon I\to M$ is called a geodesic if it satisfies $\nabla_{\dot{\gamma}}\dot{\gamma}=0$, i.e., the geodesic equation
(4.1) $\frac{d^{2}\gamma^{k}}{dt^{2}}+\sum_{i,j=0}^{n-1}\Gamma_{ij}^{k}\frac{d\gamma^{i}}{dt}\frac{d\gamma^{j}}{dt}=0,\qquad(k=0,\dots,n-1).$
Here $\Gamma^{k}_{ij}$ denote the Christoffel symbols,
(4.2) $\Gamma^{k}_{ij}=\frac{1}{2}\sum_{l=0}^{n-1}g^{kl}\bigl(\partial_{i}g_{jl}+\partial_{j}g_{il}-\partial_{l}g_{ij}\bigr),$
which are continuous functions since the metric coefficients $g_{ij}$ are $C^{1}$.
The following still holds for geodesics even in $C^{1}$-spacetimes.
Fact 4.2 ([G20, Corollary 2.4]).
In a $C^{1}$-spacetime, any geodesic does not change the causality.
In $C^{1}$-spacetime, or more generally $C^{0}$-spacetime, the curve length and Lorentz distance are defined in the usual way.
Definition 4.3 ([BEE96, Definition 4.1]).
Let $(M,g)$ be a $C^{0}$-spacetime and let $p,q\in M$. A future-directed causal curve $\gamma\colon[a,b]\to M$ means a locally Lipschitz curve whose tangent vector is future-directed and causal almost everywhere. Its $g$-length is
(4.3) $L_{g}(\gamma)\coloneqq\int_{a}^{b}\sqrt{-\,g(\dot{\gamma}(t),\dot{\gamma}(t))}\,dt,$
which is well-defined and independent of the reparametrization. Let $\Omega_{p,q}$ be the set of all future-directed causal curves from $p$ to $q$. The Lorentzian distance is defined by
(4.4) $T(p,q)\coloneqq\begin{cases}\displaystyle\sup_{\gamma\in\Omega_{p,q}}L_{g}(\gamma),&\Omega_{p,q}\neq\emptyset,\\ 0,&\Omega_{p,q}=\emptyset.\end{cases}$
Definition 4.4.
Let $(M,g)$ be a $C^{1}$-spacetime. A future-directed causal curve $\gamma\colon[a,b]\to M$ is called maximal if
$L_{g}(\gamma)\ \geq\ L_{g}(\lambda)$
for every future-directed causal curve $\lambda$ with the same endpoints as $\gamma$.
In a $C^{1}$-spacetime, the geodesic equation is merely continuous, so the geodesic ODE has locally solutions but uniqueness may fail. Indeed, [KOSS22, Example 3.2] exhibits a $C^{1}$-spacetime in which maximal timelike geodesics and null geodesics branch. Consequently, to guarantee the uniqueness of causal geodesics, we impose a non-branching condition on causal geodesics (cf. [KOSS22, Definition 3.1]). We also show that this non-branching hypothesis is actually satisfied by a class of $C^{1}$ low-regularity spacetimes treated in Appendix A; see Theorem A.3.
Definition 4.5 (cf. [KOSS22, Definition 3.1]).
Let $(M,g)$ be a $C^{1}$-spacetime. A geodesic $\gamma\colon[a,b]\to M$ is said to branch at $t_{0}\in(a,b)$ if there exists $\epsilon>0$ and a geodesic $\sigma\colon(t_{0}-\epsilon,t_{0}+\epsilon)\to M$ such that $\gamma|_{(t_{0}-\epsilon,t_{0}]}=\sigma|_{(t_{0}-\epsilon,t_{0}]}$ and $\gamma|_{(t_{0},t_{0}+\epsilon)}\cap\sigma|_{(t_{0},t_{0}+\epsilon]}=\emptyset$. $(M,g)$ is called maximally causally non-branching (MCNB) if no maximal causal geodesic branches in this sense. $(M,g)$ is called causally non-branching if no causal geodesic branches in this sense.
We define the following sufficient condition ensuring that maximal causal geodesics do not branch.
Definition 4.6.
Let $(M,g)$ be a $C^{1}$-spacetime and $p\in M$. We say that $(M,g)$ is proper at $p$ if there exists a neighborhood $U$ of $p$ such that, for any two distinct maximal causal geodesics $\gamma_{1},\gamma_{2}$ starting at $p$, the intersection set $\gamma_{1}\cap\gamma_{2}\cap U$ has no accumulation point in $U$. We say $(M,g)$ is proper if it is proper at every $p\in M$.
Lemma 4.7.
Let $(M,g)$ be a causally non-branching $C^{1}$-spacetime, proper at $p\in M$ in the sense of Definition 4.6. Suppose $\gamma_{1},\gamma_{2}:[0,1]\to M$ are maximal causal geodesics with
$\gamma_{1}(0)=\gamma_{2}(0)=p,\qquad\dot{\gamma}_{1}(0)=\dot{\gamma}_{2}(0)=v.$
Then $\gamma_{1}=\gamma_{2}$ on $[0,1]$.
Proof.
Assume $\gamma_{1}\neq\gamma_{2}$. By properness at $p$, there is $\varepsilon>0$ with $\gamma_{1}((0,\varepsilon])\cap\gamma_{2}((0,\varepsilon])=\emptyset$. Let $\gamma_{-}\colon(-\delta,0]\to M$ be a past-directed causal geodesic with $\gamma_{-}(0)=p$ and $\dot{\gamma}_{-}(0)=v$, obtained from local existence for the geodesic ODE. Define concatenations
(4.5) $\tilde{\gamma}_{i}(t)=\begin{cases}\gamma_{-}(t),&t\in(-\delta,0],\\ \gamma_{i}(t),&t\in[0,\varepsilon],\end{cases}\qquad i=1,2.$
Each $\tilde{\gamma}_{i}$ is a causal geodesic on $(-\delta,\varepsilon]$, the two coincide on $(-\delta,0]$, but they differ on $(0,\varepsilon]$. Thus we have a causal geodesic that branches at the interior point $t=0$, contradicting causal non-branching. Hence $\gamma_{1}=\gamma_{2}$ on $0,\varepsilon]$. By [KOSS22, Lemma 3.3], they coincide on the whole $[0,1]$. ∎
Since the maximal geodesic does not branch in the above situation, we can define the exponential map as follows.
Definition 4.8.
Let $(M,g)$ be a causally non-branching, proper $C^{1}$-spacetime and let $p\in M$. Then, we can define the domain of the exponential map at $p$ by
(4.6) $\mathcal{E}_{p}\coloneqq\{v\in T_{p}M;\text{$\gamma_{v}$ is defined on an interval containing $[0,1]$}\}$
and the exponential map $\exp_{p}\colon\mathcal{E}_{p}\to M$ at $p$ by
(4.7) $\exp_{p}(v)\coloneqq\gamma_{v}(1)$
where $\gamma_{v}$ is the maximal causal geodesic with $\gamma(0)=p$ and $\dot{\gamma}(0)=v$.
We adopt the definition of global hyperbolicity for $C^{1}$-spacetimes using non-totally imprisoning instead of causality.
Definition 4.9 ([S16, Definition 2.6 and Definition 3.1]).
A $C^{1}$-spacetime is called globally hyperbolic if
  1. (1)
    non-totally imprisoning, i.e., no future or past inextendible causal curve is contained in a compact set,
  2. (2)
    the causal diamond $J(p,q)\coloneqq J^{+}(p)\cap J^{-}(q)$ is compact for all $p,q\in M$ .
We show that the Lorentzian distance is strictly increasing along causal geodesics when one of its arguments is slightly shifted away from the starting point.
Lemma 4.10.
Let $(M,g)$ be a globally hyperbolic $C^{1}$-spacetime, $p,q\in M$ with $p\ll q$ and $x\colon I\to M$ be a causal geodesic starting at $q$. Then the function $T(p,x(t))$ is strictly monotonically increasing.
Proof.
Let $t_{1},t_{2}\in I$ with $t_{1}<t_{2}$. The reverse triangle inequality implies that
(4.8) $T(p,x(t_{1}))+T(x(t_{1}),x(t_{2}))\leq T(p,x(t_{2})).$
By Fact 4.2, it suffices to consider the case where $x$ is timelike or null. If $x$ is timelike, then it follows that $T(x(t_{1}),x(t_{2}))>0$, which means that $T(p,x(t))$ is strictly monotonically increasing.
If $x$ is null, assume there exist $t_{1},t_{2}\in I$ with $t_{1}<t_{2}$ such that $T(p,x(t_{1}))=T(p,x(t_{2}))$. By [G20, Proposition 2.13], there exist timelike geodesics $\gamma_{1}$ and $\gamma_{2}$ such that $L_{g}(\gamma_{1})=T(p,x(t_{1}))$ and $L_{g}(\gamma_{2})=T(p,x(t_{2}))$. Then the curve $\gamma$ obtained by concatenating $\gamma_{1}$ and $x|_{[t_{1},t_{2}]}$ is a maximal curve from $p$ to $x(t_{2})$. By [SS21, Theorem 3.3], $\gamma$ is a causal geodesic, any maximal causal curve is a causal geodesic; hence $\gamma$ is a geodesic which changes causal character along $x|_{[t_{1},t_{2}]}$, contradicting Fact 4.2. ∎
Finally, we show that condition A implies finite compactness for $C^{1}$-spacetime if the exponential map is continuous.
Theorem 4.11.
Let $(M,g)$ be a connected, globally hyperbolic, causally non-branching, and proper $C^{1}$-spacetime such that the exponential map as defined in Definition 4.8 is continuous. If $(M,g)$ satisfies Condition A, then $(M,g)$ is finitely compact.
Proof.
We prove the result by adapting [B76, Theorem 5] to our setting. By [S16, Proposition 5.9, Theorem 5.6], $(M,g)$ is strongly causal as defined for spacetimes and also causally plain by [CG12, Corollary 1.17]. Thus, by [KS18, Theorem 5.12], the 5-tuple $(M,\ll,\leq,d^{h},T)$ defined via locally Lipschitz curves (cf. [KS18, Section 5.1]) is a Lorentzian length space. The strong causality implies the non-totally imprisoning property as defined for Lorentzian length spaces (cf. [KS18, Theorem 3.26 (iv), (iii)]), and hence global hyperbolicity.
Let $p,q\in M$ with $p\ll q$ and $B>0$. From Lemma 3.2, it suffices to show that $K_{1}=\{x\in M;p\ll q\leq x,T(p,x)\leq B\}$ is compact. By Fact 2.15 and Fact 2.16 (ii), the set $K_{1}=J^{+}(q)\cap\{x\in M;T(p,x)\leq B\}$ is closed.
Let $(z_{n})_{n}$ be a sequence of points in $K_{1}$. If for every $N\in\mathbf{N}$, there exists $n\geq N$ such that $z_{n}=q$, then there is a subsequence that converges to $q\in K_{1}$. Otherwise, we may assume $z_{n}\not=q$ for all $n\in\mathbf{N}$.
By [G20, Proposition 2.13], there exists a maximal causal geodesic $G_{n}\colon[0,1]\to M$ from $q$ to $z_{n}$. Let $X_{n}\coloneqq\dot{G}_{n}(0)$, then since $\exp_{q}$ is defined, we have $G_{n}(t)=\exp_{q}(tX_{n})$. Let $L_{n}$ be the direction containing $X_{n}$ (cf. [B76, Section 2]). By the reverse triangle inequality, we have $T(p,q)\leq T(p,z_{n})\leq B$ and the set of future-pointing (causal) directions at $q$ is a compact set in the space of directions, thus, passing to a subsequence and relabeling it again by $n$, we have $L_{n}\to L_{0}\ (n\to\infty)$ where $L_{0}$ is the future-pointing direction at $q$ and $T(p,z_{n})\to b\ (n\to\infty)$ where $T(p,q)\leq b\leq B$.
Let $x(t)\colon[0,c)\to M$ be a future inextendible maximal causal geodesic starting at $q$ with $\dot{x}(0)\in L_{0}$. By Condition A, $T(p,x(t))\to\infty$ as $t\to c$, and by Lemma 4.10 the map $t\mapsto T(p,x(t))$ is strictly increasing. By Fact 2.16, $T$ is continuous and there is a unique $t_{0}\in(0,c)$ and $X_{0}\in L_{0}$ satisfying $T(p,x(t_{0}))=b$ and $\exp_{q}(X_{0})=x(t_{0})$.
Let $\lVert\ \rVert$ be the Euclidean inner product on $T_{q}M$. For each direction $L$ at $q$ let $u(L)$ be the unique future-pointing causal vector in $L$ with $\lVert u(L)\rVert=1$. Put $u_{n}\coloneqq u(L_{n})$ and $u_{0}\coloneqq u(L_{0})$. Since $L_{n}\to L_{0}\ (n\to\infty)$, we have $u_{n}\to u_{0}\ (n\to\infty)$. For each $n$ write
(4.9) $X_{n}=s_{n}u_{n},\ s_{n}\coloneqq\lVert X_{n}\rVert>0,$
and similarly $X_{0}=s_{0}u_{0}$ with $s_{0}>0$ characterized by $\exp_{q}(s_{0}u_{0})=x(t_{0})$. Define
(4.10) $F(u,s)\coloneqq T\left(p,\exp_{q}(su)\right).$
By continuity of $\exp_{q}$ and of $T$, the map $F$ is continuous in $(u,s)$ on the domain of exponential map and by Lemma 4.10, for each fixed $u$ the function $s\mapsto F(u,s)$ is strictly increasing.
We have $F(u_{n},s_{n})=T(p,z_{n})\to b\ (n\to\infty)$ and $F(u_{0},s_{0})=T(p,x(t_{0}))=b$. Suppose that $s_{n}\not\to s_{0}$. Then there exists $\varepsilon>0$ and a subsequence (still denoted $n$) with $\lvert s_{n}-s_{0}\rvert\geq\varepsilon$ for all $n$. Choose sufficiently small $\eta>0$ and take a compact set $C\coloneqq\{u\in T_{q}M\colon\lVert u-u_{0}\rVert\leq\eta,\lVert u\rVert=1\}$. Define the continuous functions on $C$ by
(4.11) $G_{+}(u)\coloneqq F(u,s_{0}+\varepsilon)-F(u,s_{0}),\qquad G_{-}(u)\coloneqq F(u,s_{0})-F(u,s_{0}-\varepsilon).$
By strict monotonicity in $s$ and the compactness of $C$, we have
(4.12) $\delta_{\varepsilon}\coloneqq\min_{u\in C}G_{+}(u)>0,\qquad\tilde{\delta}_{\varepsilon}\coloneqq\min_{u\in C}G_{-}(u)>0.$
Passing to a further subsequence, the sequence $(s_{n})_{n}$ satisfies either $s_{n}\geq s_{0}+\varepsilon$ for all $n$, or $s_{n}\leq s_{0}-\varepsilon$ for all $n$. In the first case,
(4.13) $F(u_{n},s_{n})\geq F(u_{n},s_{0}+\varepsilon)\geq F(u_{n},s_{0})+\delta_{\varepsilon},$
while in the second,
(4.14) $F(u_{n},s_{n})\ \leq F(u_{n},s_{0}-\varepsilon)\leq F(u_{n},s_{0})-\tilde{\delta}_{\varepsilon}.$
The continuity of $F$ and $u_{n}\to u_{0}$ yield $F(u_{n},s_{0})\to F(u_{0},s_{0})=b\ (n\to\infty)$, whereas we also have $F(u_{n},s_{n})\to b$ as $n\to\infty$—a contradiction. Hence $s_{n}\to s_{0}$, and with $u_{n}\to u_{0}$ we conclude $X_{n}=s_{n}u_{n}\to s_{0}u_{0}=X_{0}$
Finally, by continuity of $\exp_{q}$ we obtain $z_{n}=\exp_{q}(X_{n})\to\exp_{q}(X_{0})=x(t_{0})\in K_{1}$. Therefore every sequence in $K_{1}$ has a convergent subsequence with limit in $K_{1}$, so $K_{1}$ is compact. This proves finite compactness. ∎
The following theorem is obtained by combining Theorem 3.4, Theorem 3.8, and Theorem 4.11.
Theorem 4.12.
Let $(M,g)$ be a connected, globally hyperbolic, causally non-branching, and proper $C^{1}$-spacetime such that the exponential map as defined in Definition 4.8 is continuous. Then the following three conditions are equivalent:
  1. (1)
    $M$ is finitely compact;
  2. (2)
    $M$ is timelike Cauchy complete;
  3. (3)
    $M$ satisfies Condition A.
Appendix A Examples of $C^{1}$-Spacetimes with geodesic uniqueness
In this appendix, we introduce examples of $C^{1}$-spacetimes where the uniqueness of causal geodesics holds and the exponential map as in Definition 4.8 can be defined.
We work on $M=\mathbf{R}^{2}$ with global coordinates $(x^{0},x^{1})=(t,x)$. The metric has the form
(A.1) $g=-a(x^{0})(dx^{0})^{2}+b(x^{0})(dx^{1})^{2},$
where $a,b\in C^{1}(\mathbf{R})$ are strictly positive functions of the time coordinate $x^{0}$. A dot denotes differentiation with respect to an affine parameter $s$ along a curve, and a prime on $a,b$ denotes derivative with respect to $x^{0}$, e.g. $a^{\prime}(x^{0})=\frac{da}{dx^{0}}(x^{0})$.
Proposition A.1.
Let $g$ be as in (A.1). For any initial data
$\gamma(0)=(x^{0}_{0},x^{1}_{0}),\qquad\dot{\gamma}(0)\coloneqq(\tau_{0},\xi_{0})=(\dot{x}^{0}_{0},\dot{x}^{1}_{0})\in T_{(x^{0}_{0},x^{1}_{0})}\mathbf{R}^{2}$
with $\tau_{0}\neq 0$, there exist $\delta>0$ and a unique geodesic
$\gamma(s)=(x^{0}(s),x^{1}(s))\colon(-\delta,\delta)\to\mathbf{R}^{2}$
solving the geodesic equation with these initial conditions. In particular, every non-trivial causal geodesic satisfies $\tau_{0}\neq 0$, hence the above local uniqueness applies to all causal geodesics.
Proof.
Without loss of generality, we assume the initial data are future-directed, i.e. $\tau_{0}=\dot{x}^{0}(0)>0$. A direct computation gives the only non-trivial Christoffel symbols:
(A.2) $\Gamma^{0}_{00}=\frac{a^{\prime}}{2a},\qquad\Gamma^{0}_{11}=\frac{b^{\prime}}{2a},\qquad\Gamma^{1}_{01}=\Gamma^{1}_{10}=\frac{b^{\prime}}{2b},$
all others vanish. Thus, the geodesic equations are
(A.3) $\ddot{x}^{0}+\frac{a^{\prime}}{2a}\,(\dot{x}^{0})^{2}+\frac{b^{\prime}}{2a}\,(\dot{x}^{1})^{2}=0,\qquad\ddot{x}^{1}+\frac{b^{\prime}}{b}\,\dot{x}^{0}\,\dot{x}^{1}=0.$
The $x^{1}$-equation in (A.3) gives
(A.4) $\ddot{x}^{1}+\frac{b^{\prime}}{b}\,\dot{x}^{0}\,\dot{x}^{1}=\frac{1}{b}\frac{d}{ds}\!\bigl(b(x^{0})\,\dot{x}^{1}\bigr)=0,$
and thus, we have
(A.5) $b(x^{0})\dot{x}^{1}=\kappa\in\mathbf{R}.$
Define the constant squared speed
(A.6) $\varepsilon\coloneqq g(\dot{\gamma},\dot{\gamma})=-a(x^{0})\,(\dot{x}^{0})^{2}+b(x^{0})\,(\dot{x}^{1})^{2}.$
Combining with (A.5) yields
(A.7) $a(x^{0})\,(\dot{x}^{0})^{2}=\frac{\kappa^{2}}{b(x^{0})}-\varepsilon.$
Evaluating at $s=0$ gives $\kappa=b(x^{0}_{0})\xi_{0}$ and $\varepsilon=-a(x^{0}_{0})\tau_{0}^{2}+b(x^{0}_{0})\xi_{0}^{2}$, and hence
(A.8) $a(x^{0})(\dot{x}^{0})^{2}=\kappa^{2}\left(\frac{1}{b(x^{0})}-\frac{1}{b(x^{0}_{0})}\right)+a(x^{0}_{0})\tau_{0}^{2}.$
If $\tau_{0}\neq 0$, the left-hand side is larger than 0 at $x^{0}=x^{0}_{0}$ and stays positive near $x^{0}_{0}$ by continuity. Using the future-directed assumption $\tau_{0}>0$, we fix the positive branch
(A.9) $\dot{x}^{0}=\sqrt{\frac{\kappa^{2}/b(x^{0})-\varepsilon}{a(x^{0})}}>0.$
Separating variables yields the implicit integral
(A.10) $s-s_{0}=\int_{x^{0}_{0}}^{x^{0}(s)}\frac{\sqrt{a(u)}}{\sqrt{\kappa^{2}/b(u)-\varepsilon}}\,du.$
The integrand is continuous and strictly positive in a neighborhood of $x^{0}_{0}$, so the right-hand side is strictly increasing and continuous in $x^{0}$. By the inverse function theorem for monotone continuous functions, (A.10) determines $x^{0}$ as a unique function of $s$. From (A.5),
(A.11) $x^{1}(s)=x^{1}_{0}+\int_{s_{0}}^{s}\frac{\kappa}{b(x^{0}(u))}du,$
which is uniquely determined once $x^{0}$ is. This proves local uniqueness on some $(-\delta,\delta)$.
Finally, if a non-trivial causal geodesic satisfies $\dot{x}^{0}(0)=\tau_{0}=0$ at the initial point, the causality at $s=0$,
(A.12) $a(x^{0}_{0})\,(\dot{x}^{0}(0))^{2}+b(x^{0}_{0})\,(\dot{x}^{1}(0))^{2}\leq 0,$
forces $\dot{x}^{1}(0)=\xi_{0}=0$, a contradiction. ∎
Proposition A.2.
Under (A.1), $(\mathbf{R}^{2},g)$ is causally non-branching, and in particular maximally causally non-branching (MCNB) in the sense of Definition 4.5.
Proof.
Let $\gamma:[\alpha,\beta]\to\mathbf{R}^{2}$ be a non-trivial causal geodesic. As observed above, $\dot{x}^{0}(s)\neq 0$ for every interior point $s\in(\alpha,\beta)$. Fix such $s_{0}$. The initial data $(\gamma(s_{0}),\dot{\gamma}(s_{0}))$ then satisfy $\tau_{0}\coloneqq\dot{x}^{0}(s_{0})\neq 0$, so by Proposition A.1 the geodesic is locally unique. Hence any causal geodesic that agrees with $\gamma$ on $(s_{0}-\varepsilon,s_{0}]$ must coincide with it on $(s_{0},s_{0}+\delta)$, ruling out branching at interior points. Therefore $(\mathbf{R}^{2},g)$ is causally non-branching, and the same holds for maximal causal geodesics (MCNB). ∎
Theorem A.3.
If $\gamma\colon I\to\mathbf{R}^{2}$ and $\sigma\colon J\to\mathbf{R}^{2}$ are maximal causal geodesics for $g$ with
(A.13) $\gamma(0)=\sigma(0),\qquad\dot{\gamma}(0)=\dot{\sigma}(0)\neq 0,$
then $\gamma$ and $\sigma$ coincide on their maximal domains of definition.
Proof.
Proposition A.1 gives local agreement near $s=0$. By Proposition A.2, [KOSS22, Lemma 3.3] applies and this local agreement to coincidence on the entire maximal domains. ∎
Moreover, for the following class of $C^{1}$-spacetime which is causally non-branching and proper, the exponential map is $C^{2}$ on the future causal cone.
Theorem A.4.
Let $M=\mathbf{R}^{2}$ with coordinates $(t,x)$ and metric
$g=-a(t)\,dt^{2}+dx^{2},$
where $a\in C^{1}(\mathbf{R})$ satisfies:
  1. (H1)
    $a(t)\geq\alpha>0$ for all $t\in\mathbf{R}$,
  2. (H2)
    $\int_{t_{0}}^{+\infty}\sqrt{a(u)}\,du=+\infty$ for some $t_{0}\in\mathbf{R}$.
Then for every $q\in M$, the exponential map $\exp_{q}$ is $C^{2}$ on the cone of future-directed nonzero causal vectors in $T_{q}M$.
Proof.
Define a new time coordinate by
$\tau=T(t)\coloneqq\int_{t_{*}}^{t}\sqrt{a(u)}du,$
with fixed $t_{*}\in\mathbf{R}$. Since $a\in C^{1}$ and $a\geq\alpha>0$ by (H1), we have $\sqrt{a}\in C^{1}$ and $\sqrt{a}>0$, hence $T\in C^{2}$ is strictly increasing. Therefore $T:\mathbf{R}\to I$ is a $C^{2}$-diffeomorphism onto an open interval $I\subset\mathbf{R}$, with inverse $t=t(\tau)\in C^{2}$. By (H2), $T(t)\to+\infty$ as $t\to+\infty$, so $I$ is unbounded above: there exists $A\in[-\infty,\infty)$ such that $I=(A,\infty)$.
Let $\Phi:\mathbf{R}^{2}\to I\times\mathbf{R}$ be the $C^{2}$-diffeomorphism $\Phi(t,x)=(\tau,x)=(T(t),x)$. In the $(\tau,x)$-coordinates, using $dt=d\tau/\sqrt{a(t(\tau))}$, we compute
$g\ =\ -a(t)\,dt^{2}+dx^{2}\ =\ -a(t)\,\Bigl(\frac{d\tau}{\sqrt{a(t)}}\Bigr)^{2}+dx^{2}\ =\ -\,d\tau^{2}+dx^{2}.$
Thus $\Phi$ is an isometry from $(\mathbf{R}^{2},g)$ onto $(I\times\mathbf{R},\eta)$, where $\eta=-d\tau^{2}+dx^{2}$ is the 2-dimensional Minkowski metric.
Fix $q\in M$, and set $Q\coloneqq\Phi(q)\in I\times\mathbf{R}$. For $v\in T_{q}M$, write $w\coloneqq d\Phi_{q}(v)\in T_{Q}(I\times\mathbf{R})\simeq\mathbf{R}^{2}$. If $v$ is future-directed nonzero causal, then in Minkowski coordinates $w=(w^{\tau},w^{x})$ satisfies $w^{\tau}>0$ and $(w^{\tau})^{2}\geq(w^{x})^{2}$. Along the $\eta$-geodesic with initial data $(Q,w)$, the solution is the straight line $s\mapsto Q+s\,w$. Since $I=(A,\infty)$ is unbounded above and $w^{\tau}>0$, the entire segment $\{\tau(Q)+sw^{\tau};s\in[0,1]\}$ lies in $I$, hence this geodesic exists on $[0,1]$. Therefore the Minkowski exponential is globally defined on the future causal cone and given by
$\exp_{Q}^{\eta}(w)=Q+w.$
Since $\Phi$ is an isometry, we have
$\exp_{q}^{g}(v)\ =\ \Phi^{-1}\!\bigl(\exp_{Q}^{\eta}(d\Phi_{q}(v))\bigr)\ =\ \Phi^{-1}\!\bigl(Q+d\Phi_{q}(v)\bigr).$
Here $v\mapsto d\Phi_{q}(v)$ is smooth, $w\mapsto Q+w$ is smooth, and $\Phi^{-1}$ is $C^{2}$. Consequently the composition $v\mapsto\exp_{q}^{g}(v)$ is $C^{2}$ on the cone of future-directed nonzero causal vectors in $T_{q}M$. ∎